OpenStudy (anonymous):

I forgot how to solve logs how do i solve loga 64=3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

a^3 = 64 a = 4

OpenStudy (anonymous):

loga(a)=1 loga(b)^n=nloga(b) use both of them u'll end up getting the him's answer :)

OpenStudy (amistre64):

logs point to exponents: loga(64) = 3 points the base 'a' at the exponent '3' and lets us know it equals '64'

OpenStudy (amistre64):

loga(64) = 3 64 = a^3 a = cbrt(64)

OpenStudy (anonymous):

what about log5 (x2+7x)=log5 30....do i set them equal because the bases are the same?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

all things being the same; the differences equate

OpenStudy (amistre64):

and recall that logs only work on values greater than 0; log(-12) is undefined

OpenStudy (anonymous):

thanks i'll remember that my final is is like 15 minutes